OpenStudy (anonymous):

Differentiate x^3sin4x

OpenStudy (anonymous):

Product rule U=x^3 V=sin(4x)

OpenStudy (anonymous):

use the product rule so it will be 3x^2(sin4x) + cos(4x)(4)

OpenStudy (anonymous):

my bad its 3x^2(sin(4x) + cos(4x)(4)(x^3)